For a population that is left skewed with a mean of 25 and a standard deviation equal to 13, determine the probability of obserng ple mean of 23 or more from a sample of size 32. Click here to view page 1 of the Cumulative Standardized Normal Table What is the probability of observing a sample mean of 23 or more from a sample of size 32? P(x23) (Round to four decimal places as needed.)
Answer #1

P(Sample mean > 23)

= P(\overline{x} \geq 23)

= P(z \geq \frac{23-25}{13/\sqrt{32}})

= P(z > -0.87)

= 0.8079